# Emergence of a hump in the cubic dielectric response of glycerol: a MD study

### v1

@@ Abstract
Within this framework, the emergence of a hump in the cubic modulus can be studied as a function of the parameters $\lambda$, $\kappa$, $\alpha$ and $\gamma$. As shown in fig. A1, a hump is present if the ratio $\lambda/\kappa$ is larger than a critical value close to 0.5. The slope of the high frequency flank, controlled by $\gamma$, affects the growth of the hump with the ratio but has a negligible effect on the critical value.
Figure A2: Static cubic response parameter $\kappa$, as a function of temperature for the self and total cases. The black dashed lines correspond to $\kappa\propto 1/T$.
@@
$|X_{2,1}(0)|=\frac{1}{3}\frac{\Delta\chi_{1}(0)}{E_{0}^{2}}\frac{k_{\mathrm{B}}T}{\chi_{1}(0)^{2}\epsilon_{0}v}$ (A3)
@@
$\chi_{4}(t)=\left\langle\sum_{j=1}^{N}c_{i}(t)c_{j}(t)\right\rangle_{i}-N\langle c_{i}(t)\rangle_{i}^{2}$ (A6)
where $c(t)$ is a time correlation function chosen here to be the self dipole correlation: $c_{i}(t)=\vec{u}_{i}(t_{0})\cdot\vec{u}_{i}(t_{0}+t)$ with $\vec{u}=\vec{\mu}/\mu$. The result for each temperature is shown in fig. A3. It reaches its maximum close to the mean relaxation time and goes at long times to a plateau at $\approx 1/3$ related to the properties of the chosen correlation function. The maximum value for each temperature is shown with green stars in fig. 4a. Its increase upon cooling is a direct sign of increasingly correlated dynamics.

### v2

@@ Abstract
Within this framework, the emergence of a hump in the cubic modulus can be studied as a function of the parameters $\lambda$, $\kappa$, $\alpha$ and $\gamma$. As shown in fig. A2, a hump is present if the ratio $\lambda/\kappa$ is larger than a critical value close to 0.5. The slope of the high frequency flank, controlled by $\gamma$, affects the growth of the hump with the ratio $\lambda/\kappa$ but has a negligible effect on the critical value.
Figure A3: Static cubic response parameter $\kappa$, as a function of temperature (a) and inverse temperature (b), for the self and total cases. The black dashed lines correspond to $\kappa\propto 1/T$.
@@
Figure A4: (a) Dipole relaxation time from MD simulations and from dielectric experiments [61]. The black line corresponds to a constant activation energy $E_{\infty}$. (b) Ratio $\lambda/\kappa$, as a function of the inverse temperature, for the MD self and total cases and from experiments [20, 31, 22]. The dashed horizontal line corresponds to the threshold leading to a hump (see fig. A2).
@@
$|X_{2,1}(0)|=\frac{1}{3}\frac{\Delta_{E_{0}}\ln\chi_{1}(0)}{E_{0}^{2}}\frac{k_{\mathrm{B}}T}{\chi_{1}(0)\epsilon_{0}v}$ (A3)
@@
$\chi_{4}(t)=\frac{1}{N}\left\langle\sum_{j=1}^{N}c_{i}(t)c_{j}(t)\right\rangle_{i,t_{0}}-\langle c_{i}(t)\rangle^{2}_{i,t_{0}}$ (A6)
where $c(t)$ is a time correlation function chosen here to be the self dipole correlation: $c_{i}(t)=\vec{u}_{i}(t_{0})\cdot\vec{u}_{i}(t_{0}+t)$ with $\vec{u}=\vec{\mu}/\mu$. The result for each temperature is shown in fig. A5. It reaches its maximum close to the mean relaxation time and goes at long times to a plateau at $\approx 1/3$. This is related to the correlation function we choose, which, in the case $j=i$ in eq. A6 gives, $\langle\cos^{2}\theta\rangle=1/3$ for a random dipole orientation. This effect also manifest itself around $t=\tau_{\alpha}$ and explains the need for a vertical shift of the origin in fig. 4b.